Fermi National Laboratory


Accelerator Update

Monday May 31
The day shift began with Operations monitoring store 3549, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 2:15 PM, a power supply expert made adjustments to an AP1 transfer line Lambertson (LAM52) to improve its load regulation.

Operations terminated store 3549 at 9:23 PM.

Operators prepared the Collider Detector at Fermilab (CDF) for a twenty-minute access at 9:34 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 10:47 PM.

Operation reported at 11:11 PM that the antiproton production had been declining all shift. Operators investigated several fixes, but nothing helped. The investigation continued.

Operations put the TeV into shot setup at 11:13 PM.

Tuesday June 1
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established store 3552 at 1:38 AM with an initial luminosity of 65.44E30. Pbar resumed stacking.

At 2:13 AM, Operations had to adjust the TeV tune, chromaticity, and make a collimator move to improve the colliding beam.

At 6:50 AM, a Booster vacuum pump tripped off and caused two RF stations (BRF7 & 8) to also trip. Everything reset okay.

An expert again made adjustments to LAM52 at 10:04 AM. This work, which lasted until noon, halted stacking.

At 12:16 PM, after Operations had attempted to return to stacking, they discovered that the TeV F23 service building power had tripped off. (F23 houses the AP1 transfer line equipment.) Operations informed an expert and the Duty Electrician. They restored power and stacking resumed at 1:06 PM.

At 3:35 PM, all the Pbar Debuncher/Accumulator devices stopped triggering. Operations contacted experts who later discovered what they thought was a clock controller card problem at 4:05 PM, but it wasn't. At 4:28 PM, experts rerouted the clock signal to the card and this resolved the problem.

At 6:27 PM, a Linac RF station (LRF5) tripped off due to a water pump trip. Operations got the station back on line at 7:22 PM only to have it trip off again a few minutes later. Operator investigated and, as you can see from the pictures below, found that the water pump had tripped off and that the rectifier had developed a leak.

Operations called in a Linac water technician, who arrived at 10 PM, and began repairs.

At 11:16 PM, CDF called and asked Operations to adjust the TeV colliding beam. Operators called an expert and, per his instructions, made adjustments to the vertical tune and the longitudinal dampers. This took operators about fifteen minutes and brought the beam back into reasonable parameters.

Wednesday June 2
The midnight shift began with Operations monitoring store 3552, but with all other beam on hold due to the Linac water problem.

Linac resumed sending beam at 12:21 AM.

MiniBooNE began taking beam at 12:22 AM.

At 3 AM, CDF called complaining about the colliding beam. Operators made similar adjustments as they had before to fix the beam. This occurred again at 3:51 AM and then again at 7:15 AM.

Plans for Today
The plans for today include allowing an end-of-store study around 9 AM, allow cryo system techs to repair a wet engine flywheel at TeV sector D3, allow TeV experts to continue with other studies if a ramp can be maintained during the cryo work, and then to establish a new store this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/2/2004   email Fermilab